OpenStudy (anonymous):

Use the sum or difference formula to determine the value of the trigonometric function. sin (-pi/12)

jimthompson5910 (jim_thompson5910):

hint: pi/4 - pi/3 = -pi/12

OpenStudy (anonymous):

thnx @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910, can you help me solve this? I got \[\sqrt{2}+\sqrt{6}\div2 \]

OpenStudy (anonymous):

but i got it wrong

OpenStudy (anonymous):

I got minus not addition

jimthompson5910 (jim_thompson5910):

you should have sin(-pi/12) = sin(pi/4 - pi/3) sin(-pi/12) = sin(pi/4)cos(pi/3) - cos(pi/4)sin(pi/3) do you have that as part of your steps?

OpenStudy (anonymous):

yass

jimthompson5910 (jim_thompson5910):

so this is the full step by step picture you should have (or something similar to it) sin(-pi/12) = sin(pi/4 - pi/3) sin(-pi/12) = sin(pi/4)cos(pi/3) - cos(pi/4)sin(pi/3) sin(-pi/12) = (sqrt(2)/2)*(1/2) - (sqrt(2)/2)*(sqrt(3)/2) sin(-pi/12) = sqrt(2)/4 - sqrt(6)/4 sin(-pi/12) = (sqrt(2) - sqrt(6))/4 Therefore, \[\Large \sin\left(-\frac{\pi}{12}\right) = \frac{\sqrt{2}-\sqrt{6}}{4}\]

jimthompson5910 (jim_thompson5910):

On step 3, I got sin(pi/4) = sqrt(2)/2 sin(pi/3) = sqrt(3)/2 cos(pi/4) = sqrt(2)/2 cos(pi/3) = 1/2 using the unit circle

OpenStudy (anonymous):

omg, i forgot its 4,i got all the work but i didn't multiply the 2. thnx

jimthompson5910 (jim_thompson5910):

I gotcha, yw